OpenStudy (anonymous):

Find an equation in slope-intercept form (where possible) for the line. Through (-1, -4) and (5, -13)

OpenStudy (anonymous):

slope intercep form y = mx + b

OpenStudy (anonymous):

Use m = y2-y1/x2-1 to find your slope. m = slope. Use your linear equation: f(x) = mx + b Sub in either of those 2 points to isolate and solve for b.

OpenStudy (anonymous):

try it tish, then post if you have questions